OpenStudy (anonymous):

-x+3 y=0, x+3 y=12 ..solve

OpenStudy (anonymous):

This is probably easy for most of you..it isn't for me, sorry

ganeshie8 (ganeshie8):

add both of them

ganeshie8 (ganeshie8):

-x + 3y = 0 x + 3y =12 ------------------

ganeshie8 (ganeshie8):

solving means, isolating "x" by itself, and "y" by itself.

ganeshie8 (ganeshie8):

if u add them both, x will cancel out... disappear, and u can solve "y" first

OpenStudy (anonymous):

can you walk me through it please

ganeshie8 (ganeshie8):

sure :)

ganeshie8 (ganeshie8):

-x + 3y = 0 x + 3y =12 ------------------

ganeshie8 (ganeshie8):

add them vertically

ganeshie8 (ganeshie8):

-x + 3y = 0 x + 3y =12 ------------------ 0 + 6y = 12

ganeshie8 (ganeshie8):

see if that makes sense so far

OpenStudy (anonymous):

yes so far

ganeshie8 (ganeshie8):

so you have : 6y = 12

ganeshie8 (ganeshie8):

just divide 6 so that "y" stays all by itself

ganeshie8 (ganeshie8):

6y = 12 /6 /6 y = 2

ganeshie8 (ganeshie8):

Half part of the problem is done.

OpenStudy (anonymous):

that was really easy when you explained it what next

ganeshie8 (ganeshie8):

:) it is easy for u i can see... u still need to find "x"

OpenStudy (anonymous):

how do i find x

ganeshie8 (ganeshie8):

since u sovled "y" value already, use it

ganeshie8 (ganeshie8):

plugin y = 2 in any one of the given equations

ganeshie8 (ganeshie8):

lets plug it in second equation

ganeshie8 (ganeshie8):

x+3y=12 put y = 2, u wud get : x + 3(2) = 12

ganeshie8 (ganeshie8):

simplify it, and solve x. can u ha ? :)

OpenStudy (anonymous):

nope but im learning

ganeshie8 (ganeshie8):

good :)

ganeshie8 (ganeshie8):

x + 3(2) = 12 ^^^^

ganeshie8 (ganeshie8):

first multiply that

ganeshie8 (ganeshie8):

x + 6 = 12

ganeshie8 (ganeshie8):

next subtract 6 both sides

OpenStudy (anonymous):

6?

ganeshie8 (ganeshie8):

x + 6 = 12 - 6 -6 x = 6

ganeshie8 (ganeshie8):

Yes !!

ganeshie8 (ganeshie8):

so, ur solution is : x = 6, y = 2

OpenStudy (anonymous):

you are awesome

ganeshie8 (ganeshie8):

nope, you're awesome... u learned so fast :) keep it up lol !